OpenStudy (anonymous):

Solve the Equation: 3x^2 = 108 A) ± 6 B) ± 10 C) 6 D) 10

OpenStudy (anonymous):

Simplifying 3x^2 = 108 Solving 3x^2 = 108 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Divide each side by '3'. x^2 = 36 Simplifying x^2 = 36 Take the square root of each side: x = {-6, 6}